Saved Connect Problem CP 7-11 (algo) Fill in the missing values in the table below. Instructions: Round your answers to 1 decimal place. Nominal Gross Real Gross Year Domestic Product Domestic Product Price Index S billions) 15214.4 ($ billions) (Year 0 100) 101.7 15270.6 15604.7 15862.2 16232.2 103.8 16405.3 107.4 5 17643.1 18286.6 6 110.5 < Prev 16 of 23
Answer #1
Year Nominal GDP Real GDP Price Index
1 15214.4 14960.1 101.7
2 15850.9 15270.6 103.8
3 16405.3 15604.7 105.1
4 17036 15862.2 107.4
5
6
Price index formula=( Nominal GDP/Real GDP)*100
Year 1
Nominal GDP is $ 15214.4 billion
Price index is 101.7
Let Real GDP be x
Substituting in the formula
(15214.4/x)=101.7
101.7x=15214.4
x=15214.4/101.7=14960.1
Year 2
Real GDP is 15270.6
Price index is 103.8
Let Nominal GDP be x
Substituting in the formula
(x/15270.6)=103.8
x=15270.6 x 103.8
x=15850.9
Year 3
Price index=
(16405.3/15604.7)*100
105.1305056
Year 4
Real GDP is 15862.2
Price index is 107.4
Let Nominal GDP be x
Substituting in the formula
(x/15862.2)=107.4
x=15862.2 x 107.4
x=17036.0